OpenStudy (anonymous):

Cyndi got a 25% discount on new clothes. What decimal represents this discount? 0.25 0.025 2.5 250

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

0.25 because you divide it by 100

jimthompson5910 (jim_thompson5910):

Agreed. 25% = 25/100 = 0.25

jimthompson5910 (jim_thompson5910):

Another percent example: 63% = 63/100 = 0.63

OpenStudy (anonymous):

thank you!!!!!!@jim_thompson5910 and @mathkid911

OpenStudy (anonymous):

Your welcome